LASER-INDUCED FLUORESCENCE SPECTROSCOPY OF ZIRCONIUM MONOCARBIDE

A. J. MERER AND J. R. D. PEERS, Chemistry Department, University of British Columbia, 2036 Main Mall, Vancouver, BC, Canada V6T 1Z1; \underline{S. J. RIXON}, Department of Physics and Astronomy, University of British Columbia, 6224 Agricultural Road, Vancouver, BC, Canada V6T 1Z1.

Laser-ablated Zr metal is found to react with methane under supersonic jet-cooled conditions. New electronic bands in the 540-655 nm region have been detected by laser-induced fluorescence. Among these are bands which do not change when CD_{4} is sustituted for CH_{4} and whose lower state B value is 0.526_{5} cm^{−1}; wavelength-resolved fluorescence indicates a lower state vibrational frequency of approximately 880 cm^{−1}, suggesting that the carrier is ZrC. Analysis of the high resolution spectra is complicated by irregularities in the Zr isotope structure, indicating severe rotational perturbations. The derived bond length of r′′ = 1.73_{9} Å is consistent with those found in other 4d transition metal monocarbides.